The following data were reported in the shareholders’ equity section of the Zhang Company’s comparative balance sheets for the years ended December 31:


($ in millions)

20242023
Common stock, $ 1 par per share$ 306$ 300
Paid-in capital – excess of par174150
Retained earnings314300

During 2024, Zhang declared and paid cash dividends of $45 million. The company also declared and issued a stock dividend. No other changes occurred in shares outstanding during 2024. What was Zhang’s net income for 2024?
Answer #1

step 1

beg RE
300
change common stock6
change in pic24
cash dividend45-75 (6+24+45)
ending retained earnings
225

306-300= 6

174-150= 24


step 2) net income 314-225= 89 answer 

answered by: anonymous